OpenStudy (anonymous):

Which of these are the intercepts of y = 2x − 6? (3, 0), (0, 6) (3, 0), (0, -6) (4, 0), (0, 6) (6, 0), (0, 12)

Elsa213 (elsa213):

" 2x+y=6 To find the x-intercept, substitute in 0 for y and solve for x. 2x+(0)=6 Any expression with zero in the numerator is zero. 2x+0=6 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. 2x=6 Divide each term in the equation by 2. (2x)/(2)=(6)/(2) Simplify the left-hand side of the equation by canceling the common terms. x=(6)/(2) Simplify the right-hand side of the equation by simplifying each term. x=3 To find the y-intercept, substitute in 0 for x and solve for y. 2(0)+y=6 Multiply 2 by each term inside the parentheses. 0+y=6 Solve the equation. y=6 These are the x and y intercepts of the equation 2x+y=6. x=3, y=6 " https://answers.yahoo.com/question/index?qid=20101102113237AA1ipMa

OpenStudy (anonymous):

thank u

Elsa213 (elsa213):

Also Welcome to openstudy c: